OpenStudy (anonymous):

Express the answer in scientific notation: (1.8 x 10-2) ÷ (9 x 102) = (Place your answers in the text box as 1.234e5 where the first number is between 1 and 10 and the number after the e is the exponent for 10.)

ganeshie8 (ganeshie8):

\( (1.8 \times 10^{-2}) ÷ (9 \times 10^2) \) like this ?

OpenStudy (anonymous):

Yes sir

ganeshie8 (ganeshie8):

\(\large \frac{(1.8 \times 10^{-2}) }{(9 \times 10^2)}\)

ganeshie8 (ganeshie8):

\(\large \frac{(18 \times 10^{-3}) }{(9 \times 10^2)}\)

ganeshie8 (ganeshie8):

\(\large \frac{(2.0\times 10^{-3}) }{(10^2)}\)

ganeshie8 (ganeshie8):

you can simplify further

OpenStudy (anonymous):

the 2's?

OpenStudy (anonymous):

2.0x10^0?

ganeshie8 (ganeshie8):

\(\large 2.0 \times 10^{-3-2}\)

OpenStudy (anonymous):

-1?

ganeshie8 (ganeshie8):

\(\large 2.0 \times 10^{-5}\)

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

2.0e-5 Like that?

OpenStudy (anonymous):

(Place your answers in the text box as 1.234e5 where the first number is between 1 and 10 and the number after the e is the exponent for 10.)

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

Thanks

ganeshie8 (ganeshie8):

np :D